OpenStudy (anonymous):

The area of a garden is given by the trinomial g^2 – 2g – 24. The garden’s length is g + 4. What is the garden’s width? a)g – 2 b)g – 6 c)g – 8 d)g + 2

OpenStudy (aravindg):

this is same as telling (g+4)(g+x)=g^2 – 2g – 24

OpenStudy (aravindg):

now factorise

OpenStudy (anonymous):

but what does that mean?

OpenStudy (anonymous):

@AravindG is right

OpenStudy (aravindg):

I just used property that for rectangle area=l*b

OpenStudy (aravindg):

b=width is unknown here

OpenStudy (anonymous):

so confused

OpenStudy (anonymous):

what am I supposed to multiply all I see is letters and numbers

OpenStudy (aravindg):

area of rectangle=l*b ok?

OpenStudy (anonymous):

but what do I multiply what numbers?

OpenStudy (anonymous):

the answer is B

OpenStudy (anonymous):

100% thanks !!!!

OpenStudy (aravindg):

area=g^2 – 2g – 24 length=g + 4 width= ?

OpenStudy (aravindg):

hope that makes things cleaR

OpenStudy (raden_zaikaria):